# Revisting the Impact of Upstream Mergers with Downstream Complements and Substitutes

### v2

@@ 2 A Simple Model
In this section, I present my main results using a simple model that clarifies the underlying mechanism with maximum transparency. The model abstracts from retail and wholesale prices, as well as marginal costs, by assuming reduced-form payoffs. It also focuses on a specific scenario: the merger of two upstream suppliers of consumer goods that sell unrelated products through large retail chains offering one-stop shopping to consumers. An example of such a transaction is the merger between Procter & Gamble and Gillette in 2005 (EU Case No. COMP/M.3732). In this setting, there are only downstream complementarities arising due to one-stop-shopping economies.
In Section 3, I extend these results by accounting for firms’ pricing decisions and examining broader forms of complementarities and substitutabilities at the downstream level.
@@ 2.1 The Setting
Consumers must also incur a shopping cost $\xi\geq 0$ when visiting the retailer. This shopping cost varies among the population of consumers according to the cumulative distribution function (CDF) $G$. Final consumers observe the retailer’s product portfolio before deciding whether to incur this cost. The shopping cost could be interpreted literally as the opportunity cost of time spent in traffic and parking, selecting products, and so forth, or—taking a partial equilibrium view—as the value for the consumer of visiting its best alternative among the retailer’s competitors.
I analyze the effects of a merger between suppliers 1 and 2 under the assumption that the retailer and suppliers negotiate lump-sum carriage fees using a Nash-in-Nash bargaining protocol. I denote by $\beta\in(0,1)$ the bargaining weight of the retailer in its negotiations with each supplier, where these weights are independent of the upstream market’s ownership structure.
@@ 2.2 A Brief Discussion
Downstream Complementarities.— Complementarity in utility and gross complementarity in demand are two related but distinct notions of downstream complementarities. The first reflects that the marginal utility of a product increases when another product is present, while the second captures that the demand for one product rises when another product is available. [^7] Complementarity in utility is sometimes called “q-complementarity” or “Edgeworth-Pareto complementarity,” while gross complementarity is sometimes referred to as ”p-complementarity” (Hicks, 1970; Seidman, 1989).
@@
Why Assume a Monopoly Retailer?.— The primary aim of this paper is to challenge the prevailing view that upstream mergers weaken suppliers’ bargaining leverage with downstream complements and strengthen it with downstream substitutes. Assuming a single monopoly retailer allows me to isolate and focus on the core mechanism at play, avoiding the additional complexities introduced by competition among retailers. This approach also rules out alternative channels through which the merger could increase prices, such as the mechanism proposed by Peters (2014), as discussed in the Related Literature section. For the same reason, I assume that Nash bargaining weights remain unchanged following the merger.
@@
On Nash-in-Nash Bargaining.— In modeling upstream negotiations, I use the Nash-in-Nash bargaining protocol as my solution concept (Horn and Wolinsky, 1988; Collard-Wexler et al., 2019). This protocol has become increasingly prominent in the theoretical and empirical literature analyzing the effect of upstream negotiations on market outcomes (e.g., Dobson and Waterson, 1997; Ho and Lee, 2017). It has also become an integral part of authorities’ toolkit when evaluating upstream mergers of suppliers and vertical mergers (e.g., Farrell et al., 2011; Shapiro, 2018; Rogerson, 2020).
One potential concern with this solution concept is that it may not be suitable when suppliers’ products are strong complements for the retailer (Collard-Wexler et al., 2019). However, in the Online Appendix, I demonstrate that my main result also holds when upstream negotiations are modeled using the Shapley Value. This protocol does not have the same problems as the Nash-in-Nash solution when strong complementarities are present and can be microfounded in a noncooperative framework, as demonstrated by Stole and Zwiebel (1996) and Inderst and Wey (2003).
@@
This “all-or-nothing” negotiation approach is also consistent with practices in the US healthcare market, as noted by the Federal Trade Commission (FTC). Specifically, when evaluating hospital mergers, the FTC assumes that if an insurer fails to reach an agreement with a hospital system, it loses access to all hospitals within that system (see Farrell et al., 2011, p. 274). This assumption is likewise employed in all major academic studies estimating the effects of different policy interventions and counterfactuals in the industry (e.g., Gowrisankaran et al., 2015; Lewis and Pflum, 2015; Ho and Lee, 2017).
@@ 2.3 Upstream Mergers with Downstream Complementarities Can Increase Negotiated Prices
I now analyze how the merger of suppliers 1 and 2 affects the terms negotiated in equilibrium. In the current setting, it is straightforward to show that the merger does not affect the negotiation between the retailer and supplier 3—the supplier not involved in the transaction. Hence, without loss of generality, I focus exclusively on how the merger affects the negotiations between the retailer and suppliers 1 and 2.
Nevertheless, the inclusion or exclusion of product 3 in the retailer’s portfolio will play a crucial role in determining the merger’s impact, as will soon become apparent. I therefore compare two scenarios: one where product 3 is present ($x_{3}=1$) and one where it is absent ($x_{3}=0$), treating $x_{3}\in\{0,1\}$ as a parameter.
@@
That is, $t_{i}^{\text{pre}}$ maximizes the Nash product between the retailer’s and supplier $i$’s gains from trading with each other, assuming the retailer also reaches and agreement with the other supplier. The combined premerger payments that suppliers $1$ and $2$ can joint secure are, therefore, equal to:
@@
As noted by Shaffer (1991), Inderst and Wey (2003), and Nevo (2014), products should be substitutes for the retailer in terms of profits when they are gross substitutes in demand. This is because the incremental profits from adding two gross substitutes to the retailer’s portfolio should be smaller than the sum of the profits from adding each product individually, as the sales of one product increase when the other is not stocked.
By the same reasoning, consumer-level complementarities must also transfer upstream: when products are gross complements in demand, the retailer’s incremental profits from adding both products should be larger than the sum of the profits from adding them individually (Inderst and Wey, 2003; Easterbrook et al., 2018; Baye et al., 2018).
The problem with the conventional wisdom, however, is that it overlooks how the inclusion of a product affects the spillovers generated by other products in the retailer’s portfolio. As a result, downstream complementarities and substitutabilities do not always transfer upstream.
@@
where I am using the fact that $\Pi(x_{1},x_{2},x_{3})=\big{(}\bm{x}\cdot\bm{\pi}\big{)}\cdot G(\bm{x}\cdot\bm{v})$.
@@
However, when the retailer carries product 3 (i.e., $x_{3}=1$), an additional factor comes into play, as the right-hand side of equation (3) is no longer zero. This term—strictly positive when $G(x)$ is strictly concave and strictly negative when $G(x)$ is strictly convex—captures whether products 1 and 2 act as complements or substitutes in the profit ”spillovers” they generate for product 3, the other product in the retailer’s lineup.
@@
That is, when the increase in the retailer’s profits from product 3, achieved by adding products 1 and 2 simultaneously, is smaller than the combined profit increase from adding each product individually. In this context, this occurs when $G(x)$ is strictly concave, as it implies that adding product 1 on top of product 2 (or vice versa) does not increase consumer traffic significantly more than adding either product alone. [^8] Another way to interpret this idea is as follows: because each product can increase consumer traffic to the store, both products 1 and 2 contribute to increasing the demand for product 3. Consequently, when $G(x)$ is strictly concave, products 1 and 2 effectively act as ”substitutes” in driving the demand for product 3, making them substitutes in the positive profit spillovers they generate for this product.
@@ 2.4 Implications
Where Does the Conventional Wisdom Come From?— As evident from condition (3), downstream complementarities necessarily transfer upstream under either of the following conditions: (i) the retailer carries only the merging suppliers’ products, or (ii) $G$ is weakly convex—a common feature of Hotelling-type models with uniformly distributed consumers. In the latter case, this occurs because the merging products are (weak) complements in the profit spillovers they generate for product 3, meaning the right-hand side of (3) is weakly negative even when $x_{3}=1$.
Conditions (i) and (ii) help explain the origins of the ’conventional wisdom’ described in the previous section: Most theoretical models demonstrating that downstream complementarities (or substitutabilities) transfer upstream in the supply chain typically focus on two products (e.g., Shaffer, 1991; Inderst and Wey, 2003) and/or rely on frameworks where consumers are uniformly distributed (e.g., Baye et al., 2018). [^9] As noted in the Related Literature section, the only exception I am aware of is O’Brien and Shaffer (2005). They analyze a model where $n\geq 2$ upstream suppliers sell their products through a single monopoly retailer, with all products being gross substitutes. On page 584, they claim that because products are gross substitutes, they are substitutes for the retailer, providing a proof for that result in page 593. However, as I explain in Section 3.2, this conclusion does not hold in general; the proof in O’Brien and Shaffer (2005) contains an error., [^10] The result that downstream complementarities and substitutabilities transfer upstream also appears in Horn and Wolinsky (1988), who also consider only two products. However, their setting differs from mine as it includes strategic interaction in the downstream market, with two downstream firms facing interdependent downstream demands.
The “Holes in the Network” Theory of Harm.— Proposition 2.1 is closely related to the “Holes in the Network” (HiN) theory of harm, which was developed to explain empirical findings showing that mergers of hospitals operating in different geographic or diagnostic markets sometimes lead to higher negotiated prices (Vistnes and Sarafidis, 2013; Dafny et al., 2019).
In essence, the HiN theory states that if employers or households (the “consumers”) value having multiple hospitals included in an insurer’s network, two hospitals offering non-overlapping services—hospitals 1 and 2—can act as substitutes for the insurer in terms of profits. This occurs when consumers are willing to purchase the insurance plan if either hospital 1 or 2 is excluded but not if both are excluded (i.e., customers tolerate one but not two “holes” in the plan).
The HiN theory represents an extreme version of my argument. Indeed, the theory describes as situation where $G(v_{1}+v_{2}+v_{3})=G(v_{1}+v_{3})=G(v_{2}+v_{3})=1$ and $G(v_{3})=0$. Under these conditions, products 1 and 2 are always substitutes for the retailer in terms of profits:
@@
Intuitively, if consumers continue visiting the retailer when either product 1 or 2 is missing but stop visiting entirely when both products are absent, the two products function as ”perfect substitutes” in generating spillovers—particularly store traffic—for product 3. Furthermore, because adding product $1$ or product $2$ to the portfolio does not affect store traffic when the other is already present, the presence of product 1 does not increase the profits of product 2, and vice versa. Consequently, in the HiN theory, the substitutability of products 1 and 2 in terms of overall profits for the retailer is driven solely by their substitutability in generating spillovers for product 3.
Although the HiN theory is factually correct, its current presentation has several shortcomings, which this paper seeks to address. Chief among these is its inability to resolve the apparent contradiction with the ’conventional wisdom’ outlined earlier: How can two hospitals be complements for consumers—who value having multiple hospitals in a network—yet substitutes for insurers? This paper provides a unifying framework that resolves this contradiction and clarifies the conditions under which complementarities at the consumer level extend to the intermediary level.
Furthermore, the HiN theory was developed exclusively for cross-market mergers with common customers. As such, it focuses on a very specific type of downstream complementarity: that arising from one-stop-shopping economies. However, as I demonstrate in Section 3, the insights derived in Section 2.3 extend to other types of complementarities as well. Notably, they even apply to the case where products are gross substitutes, as I argue next.
Upstream Mergers with Downstream Substitutabilities can Reduce Negotiated Prices.— Suppose product demands take the more general form $D_{i}(x)$ for $i=1,2,3$, rather than being driven by one-stop shoppers with unit demands, as in Sections 2.12.3. Assume also that all products are gross substitutes, meaning that $D_{i}(\bm{x})$ is decreasing in $x_{j}\in\{0,1\}$ for all $i\neq j$. If the retailer generates a profit of $\pi_{i}>0$ per unit of product $i$ sold, its total profits before fixed fees are $\Pi(\bm{x})=\sum_{i}x_{i}\pi_{i}D_{i}(\bm{x})$.
@@
First, they challenge the conventional wisdom asserting that downstream substitutabilities invariably exacerbate the competitive concerns of upstream mergers, while downstream complementarities mitigate them (see, e.g., Nevo, 2014; Easterbrook et al., 2018). For downstream substitutabilities, this misconception could lead to excessively stringent cost-synergy thresholds for merger approval, potentially blocking transactions that would enhance overall welfare. For downstream complementarities, it risks prematurely classifying a merger as efficiency-driven—without further evaluation—when the merger may instead be motivated by the prospect of securing higher negotiated prices.
@@
Second, my findings call for a reassessment of how upstream conglomerate mergers are evaluated. As noted in the Introduction, the prevailing approach primarily emphasizes risks such as anticompetitive foreclosure or monopolization. For example, in its assessment of the 2005 merger between Procter & Gamble and Gillette, the European Commission approved the merger after concluding that it would not allow the parties to impose weak brands on retailers, foreclose competitors from accessing retailers’ limited shelf space, or hinder market entry of new products through bundling practices. [^11] The Commission also forced Procter & Gamble to divest its battery toothbrush business “SpinBrush,” since there were significant horizontal overlaps at the consumer level between this brand and Gillette’s “OralB” brand.
However, my results demonstrate that negotiated prices can increase through mechanisms unrelated to foreclosure, highlighting the need to revisit the evaluation criteria for these mergers.
The Effects of an Upstream Merger are Buyer-Specific.—My results provide an explanation for buyer-specific price effects following upstream mergers. For example, Thompson (2011) examines the 1998 merger between New Hanover Regional Medical Center and Columbia Cape Fear Memorial Hospital in Wilmington, North Carolina. The study documents significant variation in post-merger price changes across insurers: two insurers experienced substantial price increases, one insurer’s prices remained unaffected, and another saw a notable price decrease.
This heterogeneity can be attributed to the role of third-party products in intermediaries’ portfolios, as their presence and nature play a crucial role in determining whether downstream complementarities or substitutabilities transfer upstream. Consequently, in markets with multiple intermediaries, mergers may produce asymmetric effects—raising negotiated prices for some intermediaries while lowering them for others, depending on the specific composition of their portfolios. These buyer-specific effects also highlight the significant challenges antitrust authorities face when evaluating the overall impact of an upstream merger in markets where intermediaries carry multiple related products.
Consumer Loss Ratios to Measure Spillover Complementarities/Substitutabilities.—As evident from (3) and (6), downstream complementarities (substitutabilities) necessarily transfer upstream when the merging products are weak complements (substitutes) in the profit spillovers they generate for other items in the lineup. This naturally raises the question: how can we empirically determine whether merging products act as complements or substitutes in generating these spillovers?
@@
One potential approach—at least in principle—is to analyze consumer loss ratios, which measure the proportion of customers lost by the retailer after removing a product or set of products. Notably, consumer loss ratios have been applied in other antitrust contexts, including the evaluation of vertical mergers (e.g., U.S. v. AT&T, et al., Civil Case No. 17-2511, RJL, D.D.C. Jun. 12, 2018).
@@
For instance, in the one-stop-shopping model of Sections 2.1-2.3, note that $S_{3}(1,1)-S_{3}(0,1)>S_{3}(1,0)-S_{3}(0,0)$ if and only if:
@@
Hence, if the sum of the consumer loss ratio of products 1 and 2 is greater than the consumer loss ratio of simultaneously losing both products—i.e., if $\mathrm{CL}_{1}+\mathrm{CL}_{2}>\mathrm{CL}_{12}$—the two products are complements in generating spillovers to product $3$. [^12] At first glance, this condition might seem counterintuitive, as it indicates that products 1 and 2 are complements in spillovers when they are substitutes in terms of loss ratios. However, it is important to note that consumer loss ratios are, by definition, the negative of consumer gains. Hence, products 1 and 2 are complements in spillovers when they are complements in consumer gains.

### v4

@@ 2 A Simple Model
In this section, I present my main results using a simple model designed to clarify the underlying mechanism with maximum transparency. The model abstracts from retail and wholesale prices, as well as marginal costs, by assuming reduced-form payoffs. It focuses on a specific scenario: the merger of two upstream suppliers of consumer goods selling unrelated products through large retail chains that offer one-stop shopping to consumers. An example of such a merger is the 2005 transaction between Procter & Gamble and Gillette (EU Case No. COMP/M.3732). In this setting, downstream complementarities arise solely from one-stop-shopping economies.
In Section 3, I extend these results by incorporating firms’ pricing decisions and allowing for more general forms of complementarity and substitutability at the downstream level.
@@ 2.1 The Setting
Consumers must also incur a shopping cost $\xi\geq 0$ when visiting the retailer. This shopping cost varies among the population of consumers according to the cumulative distribution function (CDF) $G$. Final consumers observe the retailer’s product portfolio before deciding whether to incur this cost. The shopping cost can be understood literally as the opportunity cost of time spent navigating traffic, finding parking, and choosing products, or—taking a partial equilibrium view—as the value for the consumer of visiting its best alternative among the retailer’s competitors.
I analyse the effects of a merger between suppliers 1 and 2 under the assumption that the retailer and suppliers negotiate lump-sum carriage fees using a Nash-in-Nash bargaining protocol. I denote by $\beta\in(0,1)$ the bargaining weight of the retailer in its negotiations with each supplier, where these weights are independent of the upstream market’s ownership structure.
@@ 2.2 A Brief Discussion
Downstream Complementarities.— Complementarity in utility and gross complementarity in demand are two related but distinct notions of downstream complementarities. The first reflects that the marginal utility of a product increases when another product is present, while the second captures that the demand for one product rises when another product is available. [^7] Complementarity in utility is sometimes called “q-complementarity” or “Edgeworth-Pareto complementarity,” while gross complementarity is sometimes referred to as “p-complementarity” (Hicks, 1970; Seidman, 1989).
@@
Why Assume a Monopoly Retailer?.— The primary aim of this paper is to challenge the prevailing view that upstream mergers weaken suppliers’ bargaining leverage with downstream complements and strengthen it with downstream substitutes. Assuming a single monopoly retailer allows me to isolate and focus on the core mechanism at play, avoiding the additional complexities introduced by competition among retailers. This approach also rules out alternative channels through which the merger could increase prices, such as the mechanism proposed by Peters (2014) discussed in the Related Literature section. For the same reason, I assume that Nash bargaining weights remain unchanged following the merger.
@@
On Nash-in-Nash Bargaining.— I model upstream negotiations using the Nash-in-Nash bargaining protocol (Horn and Wolinsky, 1988; Collard-Wexler et al., 2019). This protocol has become increasingly prominent in the theoretical and empirical literature analyzing the effect of upstream negotiations on market outcomes (e.g., Dobson and Waterson, 1997; Ho and Lee, 2017). It has also become an integral part of authorities’ toolkit when evaluating upstream mergers of suppliers and vertical mergers (e.g., Farrell et al., 2011; Shapiro, 2018; Rogerson, 2020).
One potential concern with this solution concept is that it may not be suitable when suppliers’ products are strong complements for the retailer (Collard-Wexler et al., 2019). However, in the Online Appendix, I show that my main result also holds when upstream negotiations are modeled using the Shapley Value. This protocol does not have the same problems as the Nash-in-Nash solution when strong complementarities are present and can be microfounded in a noncooperative framework, as demonstrated by Stole and Zwiebel (1996) and Inderst and Wey (2003).
@@
This “all-or-nothing” negotiation approach is also commonly observed in the U.S. healthcare market. When evaluating hospital mergers, the Federal Trade Commission (FTC) accounts for the fact that insurers often lose access to an entire hospital system if negotiations break down with any one hospital within it (see Farrell et al., 2011, p. 274). Consistent with this practice, major academic studies estimating the effects of policy interventions and counterfactuals in the industry also model bargaining on an all-or-nothing basis (e.g., Gowrisankaran et al., 2015; Lewis and Pflum, 2015; Ho and Lee, 2017).
@@ 2.3 Upstream Mergers with Downstream Complementarities Can Increase Negotiated Prices
I now analyse how the merger of suppliers 1 and 2 affects the terms negotiated in equilibrium. In the current setting, it is straightforward to show that the merger does not affect the negotiation between the retailer and supplier 3—the supplier not involved in the transaction. Hence, without loss of generality, I focus exclusively on how the merger affects the negotiations between the retailer and suppliers 1 and 2.
However, the inclusion or exclusion of product 3 in the retailer’s portfolio will play a crucial role in determining the merger’s impact, as will soon become apparent. I therefore compare two scenarios: one where product 3 is present ($x_{3}=1$) and one where it is absent ($x_{3}=0$), treating $x_{3}\in\{0,1\}$ as a parameter.
@@
That is, $t_{i}^{\text{pre}}$ maximises the Nash product between the retailer’s and supplier $i$’s gains from trading with each other, assuming the retailer also reaches an agreement with the other supplier. The combined premerger payments that suppliers $1$ and $2$ can jointly secure are, therefore, equal to:
@@
As noted by Shaffer (1991), Inderst and Wey (2003), and O’Brien and Shaffer (2005), products that are gross substitutes in demand should also be substitutes in profits from the retailer’s perspective. The reason is that when two such products are added to the retailer’s portfolio, some consumers switch between them rather than generating entirely new sales. As a result, the incremental profit from offering both products is less than the sum of the profits from offering each one in isolation.
By the same logic, consumer-level complementarities should also carry over to the retailer: when products are gross complements in demand, offering both increases the likelihood of joint purchases, making the combined profit greater than the sum of the stand-alone profits (Inderst and Wey, 2003; Baye et al., 2018).
The problem with the conventional wisdom, however, is that it overlooks how the inclusion of a product affects the spillovers created for other products in the retailer’s portfolio. As a result, downstream complementarities and substitutabilities do not always transfer upstream.
@@
where I use the fact that $\Pi(x_{1},x_{2},x_{3})=\big{(}\bm{x}\cdot\bm{\pi}\big{)}\cdot G(\bm{x}\cdot\bm{v})$.
@@
However, when the retailer carries product 3 (i.e., $x_{3}=1$), an additional factor comes into play, as the right-hand side of equation (3) is no longer zero. This term—strictly positive when $G(x)$ is strictly concave and strictly negative when $G(x)$ is strictly convex—captures whether products 1 and 2 act as complements or substitutes in the profit “spillovers” they generate for product 3, the other product in the retailer’s lineup.
@@
That is, when the increase in the retailer’s profits from product 3, achieved by adding products 1 and 2 simultaneously, is smaller than the combined profit increase from adding each product individually. In this context, this occurs when $G(x)$ is strictly concave, as it implies that adding product 1 on top of product 2 (or vice versa) does not increase consumer traffic significantly more than adding either product alone.
@@ 2.4 Implications
Where Does the Conventional Wisdom Come From?— As evident from condition (3), downstream complementarities necessarily transfer upstream when the retailer carries only the merging suppliers’ products. This fact can help explain the origins of the conventional wisdom described in the previous section. Indeed, as discussed in the Related Literature section, most theoretical models showing that downstream complementarities (or substitutabilities) transfer upstream typically either focus on two products (e.g., Shaffer, 1991; Inderst and Wey, 2003; Baye et al., 2018). [^8] The result that downstream complementarities and substitutabilities transfer upstream also appears in Horn and Wolinsky (1988), who also consider only two products. However, their setting differs from mine as it includes strategic interaction in the downstream market, with two downstream firms facing interdependent downstream demands.
The “Holes in the Network” Theory of Harm.— Proposition 2.1 is closely related to the “Holes in the Network” (HiN) theory of harm, developed to explain why mergers of hospitals operating in different geographic or diagnostic markets can sometimes raise negotiated prices (Vistnes and Sarafidis, 2013; Dafny et al., 2019).
In essence, the HiN theory states that if employers or households (the “consumers”) value having multiple hospitals included in an insurer’s network, two hospitals offering non-overlapping services—hospitals 1 and 2—can act as substitutes for the insurer in terms of profits. This occurs when consumers are willing to purchase the insurance plan if either hospital 1 or 2 is excluded, but not if both are excluded (i.e., customers tolerate one but not two “holes” in the plan).
The HiN theory represents an extreme version of my argument. Indeed, the theory describes a situation where $G(v_{1}+v_{2}+v_{3})=G(v_{1}+v_{3})=G(v_{2}+v_{3})=1$ and $G(v_{3})=0$. Under these conditions, products 1 and 2 are always substitutes for the retailer in terms of profits:
@@
Intuitively, if consumers continue visiting the retailer when either product 1 or 2 is missing but stop visiting entirely when both products are absent, the two products function as perfect substitutes in generating spillovers—particularly store traffic—for product 3. Furthermore, because adding product $1$ or product $2$ to the portfolio does not affect store traffic when the other is already present, the presence of product 1 does not increase the profits of product 2, and vice versa. Consequently, in the HiN theory, the substitutability of products 1 and 2 in terms of profits for the retailer is driven solely by their substitutability in generating spillovers for product 3.
Although the HiN theory is insightful, its current presentation has several shortcomings, which this paper addresses. The main one is its inability to resolve the apparent contradiction with the conventional wisdom outlined earlier: How can two hospitals be complements for consumers—who value broad networks—yet substitutes for insurers? This paper provides a unifying framework that resolves this contradiction and clarifies the conditions under which complementarities at the consumer level extend to the intermediary level.
Moreover, the HiN theory focuses narrowly on downstream complementarities driven by one-stop-shopping economies—that is, the desire of employers and households to purchase plans covering a wide range of health needs. As I demonstrate in Section 3, however, the insights extend to other types of complementarities as well. Notably, they even apply when products are gross substitutes, as I argue next.
Upstream Mergers with Downstream Substitutabilities can Reduce Negotiated Prices.— Suppose product demands take the more general form $D_{i}(\bm{x})$ for $i=1,2,3$, rather than being driven by one-stop shoppers with unit demands, as in Sections 2.12.3. Assume also that all products are gross substitutes, meaning that $D_{i}(\bm{x})$ is decreasing in $x_{j}\in\{0,1\}$ for all $i\neq j$. If the retailer generates a profit of $\pi_{i}>0$ per unit of product $i$ sold, its total profits before fixed fees are $\Pi(\bm{x})=\sum_{i}x_{i}\pi_{i}D_{i}(\bm{x})$.
@@
The Effects of an Upstream Merger can be Buyer-Specific.— My results help explain why upstream mergers can produce buyer-specific price effects. For instance, Thompson (2011) analyses the 1998 horizontal merger between New Hanover Regional Medical Center and Columbia Cape Fear Memorial Hospital in Wilmington, North Carolina. The study documents significant variation in post-merger price changes across insurers (also referred to as “managed care organisations” or MCOs): two insurers faced substantial price increases, one insurer’s prices remained stable, and another experienced notable price decreases.
Similarly, Haas-Wilson and Garmon (2009) examine two horizontal hospital mergers in Chicago, Illinois, in 2000: the acquisition of Highland Park Hospital (HPH) by Evanston Northwestern Healthcare (ENH), and the merger between Provena St. Therese Medical Center (STMC) and Victory Memorial Hospital (VMH). After the ENH/HPH merger, four of the five active insurers in the market experienced statistically significant price increases, while the fifth avoided price increases altogether. Following the STMC/VMH merger, three of the same five insurers saw statistically significant price decreases, one insurer experienced no significant change, and one faced significant price increases.
Notably, the insurer that suffered the largest price increases after the ENH/HPH merger was also the only one to experience price increases following the STMC/VMH merger. This leads Haas-Wilson and Garmon (2009, p. 20) to conclude:
One interesting aspect of the price change estimates from the MCO data is their variance across MCOs, suggesting that some MCOs may be more vulnerable to hospital mergers than others.
This observed variation aligns with the portfolio mechanism proposed here: differences in insurers’ networks likely determine whether merging hospitals act as complements or substitutes for each insurer. However, neither study provides detailed insurer-level data—to protect insurers’ anonymity—and therefore cannot directly confirm this mechanism. [^9] If insurer characteristics were available, we could test for portfolio effects in two ways. First, insurers with smaller networks should be more likely to experience price increases, since the merging hospitals represent a larger share of their network, limiting the potential for portfolio spillover effects. Second, as I discuss below, we could examine consumer loss ratios to test whether insurers experiencing lower post-merger prices were those for whom the merging hospitals functioned as complements in generating spillover effects—a necessary condition for the hospitals to be substitutes from the perspective of consumers but complements from the perspective of insurers. Nevertheless, alternative explanations struggle to account for these patterns, especially in the New Hanover/Cape Fear and STMC/VMH mergers.
Indeed, both mergers were horizontal, involving hospitals that competed—at least to some extent—in the same geographic area and diagnostic market (Thompson, 2011; Haas-Wilson and Garmon, 2009). Thus, the conventional view would predict price increases across all insurers, not insurer-specific price decreases. A similar prediction follows from Lewis and Pflum’s (2015) argument that mergers enhance hospitals’ bargaining skills, and from Peters’s (2014) “recapture effect,” as mergers always increase recapture rates (see the Related Literature Section for an overview of these alternative mechanisms).
Hence, in the absence of portfolio effects, explaining why some insurers experienced price increases while others saw price decreases would require combining one or more of these theories with insurer-specific synergies. However, it is unclear why a hospital system would realize synergies with certain insurers but not others. A more straightforward explanation is that these buyer-specific effects are a natural consequence of the portfolio mechanism developed in this paper.
@@
First, they challenge the conventional wisdom asserting that downstream substitutabilities invariably exacerbate the competitive concerns of upstream mergers, while downstream complementarities mitigate them. For downstream substitutabilities, this misconception could lead to excessively stringent cost-synergy thresholds for merger approval, potentially blocking transactions that would enhance overall welfare. For downstream complementarities, it risks prematurely classifying a merger as efficiency-driven—without further evaluation—when the merger may instead be motivated by the prospect of securing higher negotiated prices.
@@
Second, my findings call for a reassessment of how upstream conglomerate mergers are evaluated. As noted in the Introduction, the prevailing approach primarily emphasizes risks such as anticompetitive foreclosure or monopolisation. For example, in its assessment of the 2005 merger between Procter & Gamble and Gillette, the European Commission approved the merger after concluding that it would not allow the parties to impose weak brands on retailers, foreclose competitors from accessing retailers’ limited shelf space, or hinder market entry of new products through bundling practices. [^10] The Commission also forced Procter & Gamble to divest its battery toothbrush business “SpinBrush,” since there were significant horizontal overlaps at the consumer level between this brand and Gillette’s “OralB” brand. However, my results demonstrate that negotiated prices can increase through mechanisms unrelated to foreclosure, highlighting the need to revisit the evaluation criteria for these mergers.
Third, the fact that upstream mergers can raise negotiated prices for some intermediaries while lowering them for others implies that such mergers may have complex distributional consequences. Some intermediaries and their consumers may benefit, while others may be harmed. Authorities should therefore recognize that merger effects need not be uniform across buyers—and that accurately weighing these gains and losses typically requires a structural model.
Because estimating such models can be costly, it would be useful to first assess whether profit spillovers are likely to cause the merging products to act as complements for some intermediaries and substitutes for others. As I explain next, consumer loss ratios—sometimes estimable without a full structural model, using data from blackouts, delistings, or consumer surveys—can offer a practical first approximation for this purpose.
@@
Consumer Loss Ratios to Measure Spillover Complementarities/Substitutabilities.—My results suggests that analyzing consumer loss ratios—which measure the proportion of customers lost by the retailer after removing a product or set of products—can provide a rough approximation as to whether the merging products are complements or substitutes in spillovers for a given intermediary. Notably, consumer loss ratios have been applied in other antitrust contexts, including the evaluation of vertical mergers (e.g., U.S. v. AT&T, et al., Civil Case No. 17-2511, RJL, D.D.C. Jun. 12, 2018).
@@
Indeed, in the one-stop-shopping model of Sections 2.1-2.3, note that $S_{3}(1,1)-S_{3}(0,1)>S_{3}(1,0)-S_{3}(0,0)$ if and only if:
@@
Hence, if the sum of the consumer loss ratio of products 1 and 2 is greater than the consumer loss ratio of simultaneously losing both products—i.e., if $\mathrm{CL}_{1}+\mathrm{CL}_{2}>\mathrm{CL}_{12}$—the two products are complements in generating spillovers to product $3$.